OpenStudy (anonymous):

can someone show the solution to the integral of sin^3 4x cos^5 4x dx

OpenStudy (amistre64):

cos^5 can be converted into sin^3*cos^?

OpenStudy (turingtest):

strip out a sin(4x) and you get\[\cos^5(4x)\sin^2(4x)\sin(4x)\]now use\[\sin^2u=1-\cos^2u\]and then a simple -sub

OpenStudy (turingtest):

u-sub

OpenStudy (turingtest):

amistre doesn't like my answer :(

OpenStudy (anonymous):

does my u=cos 4x?

OpenStudy (amistre64):

\[ sin^3(4x) cos^5(4x)\] \[ sin^3(4x) (cos^2(4x))^2cos(4x)\] \[ sin^3(4x)(1-sin^2(4x))^2cos^3(4x)\] \[ sin^3(4x)(1-2sin^2(4x)+sin^4(4x))cos^3(4x)\] \[ (sin^3(4x)-2sin^5(4x)+sin^7(4x))cos^3(4x)\] and that is just alot of int f(x)f'(x) stuff

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

if you do it my way amistre's is clearly different

OpenStudy (amistre64):

i undid the cos part, i think you undid the sin?

OpenStudy (anonymous):

yes the rule said if both odd its either way

OpenStudy (amistre64):

and i got a ^3 clingon lol

OpenStudy (turingtest):

yeah\[\cos^5(4x)\sin^2(4x)\sin(4x)=\cos^5(4x)(1-\cos^2)\sin(4x)\]\[=\cos^5(4x)\sin(4x)-\cos^7(4x)\sin(4x)\]at least that's how mine unfolds...

OpenStudy (amistre64):

i spose the smaller odd is a little easier to play with

OpenStudy (turingtest):

\[u=\cos(4x)\implies du=-4\sin(4x)dx\]

OpenStudy (turingtest):

just a little maybe....

OpenStudy (turingtest):

always good to see more than one way though

OpenStudy (anonymous):

my final answer is -cos^6 4x/24 + cos^8 4x/56 + c...please correct my mistakes..

OpenStudy (turingtest):

not sure about the 56

OpenStudy (anonymous):

yeah its 32, right?

OpenStudy (turingtest):

after the sub we have a 1/4 in front of it, right? than we divide by the exponent, which is...?

OpenStudy (turingtest):

\[u=\cos(4x)\implies du=-4\sin(4x)dx\implies-\frac14du=\sin(4x)dx\]

OpenStudy (turingtest):

oh yeah, that's 32, sorry I missed that comment

OpenStudy (anonymous):

thanks turning test..where you from?

OpenStudy (turingtest):

I am from the United States but live in Mexico you?

OpenStudy (anonymous):

Philippines..i'm studying bse math..

OpenStudy (turingtest):

Cool well, you're welcome for the help and I hope to see you around!